Discrete Mathematics (Mathematics 60610), Spring Semester 2008 January 3, 2008

Instructor: Andrew Sommese (sommese@nd.edu)

291 Hurley Phone: 631-6498

Time/Place: MWF: 10:40 - 11:30; DeBartolo 244

Class Website: www.nd.edu/~sommese/math08S610

Office Hours: I am available after class and 12:30-1:30 Wednesdays in my office.

Examinations, homework, and grades: There will be two examinations worth 100 points and a final examination worth 150 points. One or both of the two nonfinal examinations will be take home. The final exam will be a two hour exam covering all the material of the course with emphasis on the material covered after the second exam.

Homework will be assigned regularly, and is an integral part of the course. Typically homework will be assigned on Wednesday and collected the following Wednesday. You are allowed and encouraged to use your notes and any library books while doing the homework.

Both examinations and the homework are conducted under the honor code. While cooperation in doing homework is permitted (and encouraged), copying is not.

Homework will be worth 100 points. Thus the total number of possible points for the semester is 450. The numerical break points for letter grades (A, A-, B+,…) will be based only on the test scores and the homework.

We will follow the paperback book P. Cameron, “Combinatorics: Topics, Techniques, Algorithms,” Cambridge University Press (1995), which is available at the Notre Dame Bookstore. Implementation of algorithms is of basic importance in discrete mathematics (and a lot of fun also). We will mainly use Maple for this purpose (you may use Mathematica as you prefer).

Exam 1: Monday, February 18 probably take home. Exam 2: Monday, April 7 probably take home. Final: time and place will be announced.

The most recent version of this handout plus other useful materials can be found

at www.nd.edu/~sommese/math08S610.